### v3

@@ D.2 VQA
[H] Harvard-FairVLMed10k VQA performance - Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] Harvard-FairVLMed10k VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] Harvard-FairVLMed10k VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] Harvard-FairVLMed10k VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] Harvard-FairVLMed10k VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] PathVQA VQA performance — Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] PathVQA VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] PathVQA VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] PathVQA VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] PathVQA VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] SLAKE VQA performance — Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] SLAKE VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] SLAKE VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] SLAKE VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] SLAKE VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] VQA-RAD VQA performance — Qwen & Lingshu family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric Lingshu Qwen25-VL Qwen2-VL Lingshu Qwen25-VL Qwen2-VL Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] VQA-RAD VQA performance — Gemma family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric MedGemma Gemma3 MedGemma Gemma3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] VQA-RAD VQA performance — LLaVA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric LLaVA-Med LLaVA-1.5 LLaVA-Med LLaVA-1.5 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] VQA-RAD VQA performance — VILA family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric VILA-M3 VILA VILA-M3 VILA Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score
[H] VQA-RAD VQA performance — InternVL3 family. All values are shown as percentages. Each entry is the mean score, with the subscript showing the standard deviation, and parentheses indicating the confidence interval when available.
Off-the-shelf Fine-tuned (ML) Metric InternVL3 InternVL3 Open-ended BLEU-1 ROUGE-L Exact Match Accuracy F1 Score GPT Score Closed-ended Accuracy F1 Score Overall Exact Match F1 Score GPT Score

### v4

@@ D.2 VQA
| Off-the-shelf | Fine-tuned (ML) | |||||
| Metric | Lingshu | Qwen25-VL | Qwen2-VL | Lingshu | Qwen25-VL | Qwen2-VL |
| Open-ended | ||||||
| BLEU-1 | 10.730.21 [-1pt](10.33, 11.17) | 4.920.07 [-1pt](4.79, 5.05) | 11.940.32 [-1pt](11.27, 12.53) | 36.510.50 [-1pt](35.55, 37.53) | 43.150.52 [-1pt](42.09, 44.12) | 44.330.53 [-1pt](43.28, 45.39) |
| ROUGE-L | 14.620.24 [-1pt](14.15, 15.11) | 7.340.09 [-1pt](7.16, 7.53) | 15.160.34 [-1pt](14.47, 15.80) | 38.280.52 [-1pt](37.29, 39.32) | 45.190.52 [-1pt](44.13, 46.14) | 46.900.52 [-1pt](45.84, 47.92) |
| Exact Match | 10.680.22 [-1pt](10.26, 11.13) | 4.830.06 [-1pt](4.70, 4.95) | 12.390.35 [-1pt](11.64, 13.08) | 41.470.56 [-1pt](40.43, 42.54) | 49.190.55 [-1pt](48.09, 50.24) | 51.800.55 [-1pt](50.69, 52.83) |
| Accuracy | 33.821.06 [-1pt](31.70, 35.88) | 19.720.92 [-1pt](17.96, 21.47) | 21.430.92 [-1pt](19.64, 23.20) | 12.820.73 [-1pt](11.43, 14.26) | 17.640.83 [-1pt](16.07, 19.32) | 16.590.87 [-1pt](15.02, 18.33) |
| F1 Score | 24.050.28 [-1pt](23.54, 24.60) | 15.660.21 [-1pt](15.26, 16.08) | 23.520.36 [-1pt](22.77, 24.18) | 42.180.53 [-1pt](41.19, 43.22) | 48.660.53 [-1pt](47.60, 49.64) | 50.330.53 [-1pt](49.24, 51.39) |
| GPT Score | 60.080.63 [-1pt](58.88, 61.30) | 48.400.50 [-1pt](47.43, 49.34) | 50.590.57 [-1pt](49.45, 51.66) | 57.260.62 [-1pt](56.04, 58.44) | 60.230.59 [-1pt](59.12, 61.39) | 58.730.64 [-1pt](57.52, 59.92) |
| Closed-ended | ||||||
| Accuracy | 66.151.07 [-1pt](64.10, 68.12) | 61.431.07 [-1pt](59.39, 63.57) | 58.591.11 [-1pt](56.45, 60.82) | 73.910.97 [-1pt](72.10, 75.83) | 85.000.80 [-1pt](83.34, 86.52) | 85.370.79 [-1pt](83.81, 86.85) |
| F1 Score | 66.151.07 [-1pt](64.10, 68.12) | 16.070.70 [-1pt](14.65, 17.50) | 58.591.11 [-1pt](56.45, 60.82) | 73.910.97 [-1pt](72.10, 75.83) | 85.000.80 [-1pt](83.34, 86.52) | 85.370.79 [-1pt](83.81, 86.85) |
| Overall | ||||||
| Exact Match | 38.400.72 [-1pt](37.04, 39.80) | 9.300.36 [-1pt](8.62, 10.02) | 35.480.72 [-1pt](34.09, 36.87) | 57.680.62 [-1pt](56.54, 58.88) | 67.080.57 [-1pt](65.94, 68.18) | 68.580.55 [-1pt](67.45, 69.63) |
| F1 Score | 45.090.66 [-1pt](43.86, 46.36) | 15.870.37 [-1pt](15.13, 16.61) | 41.050.67 [-1pt](39.73, 42.28) | 58.030.61 [-1pt](56.90, 59.22) | 66.810.57 [-1pt](65.65, 67.90) | 67.840.55 [-1pt](66.71, 68.89) |
| GPT Score | 63.110.64 [-1pt](61.89, 64.38) | 54.910.60 [-1pt](53.76, 56.07) | 54.590.64 [-1pt](53.30, 55.87) | 65.580.58 [-1pt](64.44, 66.74) | 72.610.54 [-1pt](71.50, 73.69) | 72.040.55 [-1pt](70.93, 73.09) |
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | MedGemma | Gemma3 | MedGemma | Gemma3 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | LLaVA-Med | LLaVA-1.5 | LLaVA-Med | LLaVA-1.5 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | VILA-M3 | VILA | VILA-M3 | VILA |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |
| Metric | InternVL3 | InternVL3 |
| Open-ended | ||
| BLEU-1 | ||
| ROUGE-L | ||
| Exact Match | ||
| Accuracy | ||
| F1 Score | ||
| GPT Score | ||
| Closed-ended | ||
| Accuracy | ||
| F1 Score | ||
| Overall | ||
| Exact Match | ||
| F1 Score | ||
| GPT Score | ||
| Off-the-shelf | Fine-tuned (ML) | |||||
| Metric | Lingshu | Qwen25-VL | Qwen2-VL | Lingshu | Qwen25-VL | Qwen2-VL |
| Open-ended | ||||||
| BLEU-1 | 3.080.24 [-1pt](2.64, 3.56) | 0.590.02 [-1pt](0.55, 0.64) | 0.770.03 [-1pt](0.71, 0.83) | 6.980.42 [-1pt](6.18, 7.82) | 35.880.80 [-1pt](34.41, 37.41) | 36.600.84 [-1pt](34.97, 38.19) |
| ROUGE-L | 3.830.24 [-1pt](3.37, 4.31) | 1.020.04 [-1pt](0.95, 1.09) | 1.300.05 [-1pt](1.22, 1.40) | 7.370.43 [-1pt](6.56, 8.20) | 37.250.79 [-1pt](35.75, 38.79) | 37.980.83 [-1pt](36.33, 39.60) |
| Exact Match | 3.130.24 [-1pt](2.69, 3.61) | 0.580.02 [-1pt](0.54, 0.62) | 0.750.03 [-1pt](0.70, 0.81) | 7.780.44 [-1pt](6.94, 8.62) | 38.490.80 [-1pt](36.98, 40.01) | 39.440.85 [-1pt](37.75, 41.07) |
| Accuracy | 17.610.63 [-1pt](16.43, 18.81) | 10.320.52 [-1pt](9.30, 11.37) | 11.540.54 [-1pt](10.49, 12.61) | 6.240.42 [-1pt](5.45, 7.09) | 34.190.82 [-1pt](32.70, 35.79) | 34.450.85 [-1pt](32.85, 36.03) |
| F1 Score | 5.260.25 [-1pt](4.79, 5.77) | 2.220.08 [-1pt](2.06, 2.39) | 2.720.10 [-1pt](2.53, 2.92) | 7.390.43 [-1pt](6.58, 8.24) | 37.340.80 [-1pt](35.84, 38.89) | 38.060.83 [-1pt](36.43, 39.68) |
| GPT Score | 44.590.49 [-1pt](43.62, 45.54) | 40.760.40 [-1pt](39.98, 41.57) | 41.660.42 [-1pt](40.81, 42.47) | 13.190.46 [-1pt](12.32, 14.13) | 50.210.66 [-1pt](48.97, 51.50) | 51.210.67 [-1pt](49.84, 52.52) |
| Closed-ended | ||||||
| Accuracy | 81.600.66 [-1pt](80.33, 82.92) | 62.360.82 [-1pt](60.76, 63.92) | 68.250.82 [-1pt](66.67, 69.89) | 63.460.82 [-1pt](61.82, 65.02) | 92.350.46 [-1pt](91.39, 93.20) | 92.550.44 [-1pt](91.62, 93.41) |
| F1 Score | 81.600.66 [-1pt](80.33, 82.92) | 38.980.82 [-1pt](37.37, 40.63) | 68.250.82 [-1pt](66.67, 69.89) | 63.460.82 [-1pt](61.82, 65.02) | 92.350.46 [-1pt](91.39, 93.20) | 92.550.44 [-1pt](91.62, 93.41) |
| Overall | ||||||
| Exact Match | 42.410.60 [-1pt](41.25, 43.59) | 19.450.48 [-1pt](18.51, 20.40) | 34.540.58 [-1pt](33.37, 35.69) | 35.650.58 [-1pt](34.53, 36.73) | 65.450.57 [-1pt](64.30, 66.56) | 66.030.59 [-1pt](64.90, 67.16) |
| F1 Score | 43.470.59 [-1pt](42.35, 44.63) | 20.620.48 [-1pt](19.68, 21.55) | 35.520.57 [-1pt](34.37, 36.65) | 35.460.58 [-1pt](34.32, 36.54) | 64.880.57 [-1pt](63.77, 66.01) | 65.330.59 [-1pt](64.20, 66.46) |
| GPT Score | 63.120.48 [-1pt](62.16, 64.04) | 51.570.47 [-1pt](50.68, 52.48) | 54.970.49 [-1pt](54.03, 55.90) | 38.350.57 [-1pt](37.25, 39.43) | 71.300.48 [-1pt](70.36, 72.32) | 71.900.49 [-1pt](70.92, 72.84) |
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | MedGemma | Gemma3 | MedGemma | Gemma3 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | LLaVA-Med | LLaVA-1.5 | LLaVA-Med | LLaVA-1.5 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | VILA-M3 | VILA | VILA-M3 | VILA |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |
| Metric | InternVL3 | InternVL3 |
| Open-ended | ||
| BLEU-1 | ||
| ROUGE-L | ||
| Exact Match | ||
| Accuracy | ||
| F1 Score | ||
| GPT Score | ||
| Closed-ended | ||
| Accuracy | ||
| F1 Score | ||
| Overall | ||
| Exact Match | ||
| F1 Score | ||
| GPT Score | ||
| Off-the-shelf | Fine-tuned (ML) | |||||
| Metric | Lingshu | Qwen25-VL | Qwen2-VL | Lingshu | Qwen25-VL | Qwen2-VL |
| Open-ended | ||||||
| BLEU-1 | 13.441.08 [-1pt](11.37, 15.65) | 1.170.05 [-1pt](1.08, 1.27) | 4.510.31 [-1pt](3.95, 5.15) | 58.851.76 [-1pt](55.44, 62.15) | 82.831.37 [-1pt](79.96, 85.50) | 82.011.45 [-1pt](79.26, 84.70) |
| ROUGE-L | 15.901.10 [-1pt](13.86, 18.07) | 2.250.09 [-1pt](2.07, 2.41) | 7.470.45 [-1pt](6.64, 8.39) | 60.581.73 [-1pt](57.26, 64.08) | 83.301.34 [-1pt](80.59, 85.97) | 82.231.43 [-1pt](79.53, 84.85) |
| Exact Match | 13.431.09 [-1pt](11.32, 15.61) | 1.160.05 [-1pt](1.06, 1.25) | 4.490.31 [-1pt](3.93, 5.12) | 60.301.74 [-1pt](56.85, 63.75) | 83.221.33 [-1pt](80.61, 85.86) | 82.601.41 [-1pt](79.90, 85.21) |
| Accuracy | 54.301.95 [-1pt](50.38, 57.95) | 43.981.94 [-1pt](40.16, 47.77) | 45.191.91 [-1pt](41.33, 48.82) | 63.981.91 [-1pt](60.31, 67.81) | 81.921.51 [-1pt](78.88, 84.92) | 80.091.63 [-1pt](76.75, 83.12) |
| F1 Score | 17.861.11 [-1pt](15.77, 20.11) | 4.230.19 [-1pt](3.84, 4.60) | 9.300.48 [-1pt](8.38, 10.25) | 61.791.72 [-1pt](58.49, 65.05) | 83.941.33 [-1pt](81.24, 86.55) | 83.021.42 [-1pt](80.32, 85.72) |
| GPT Score | 71.971.20 [-1pt](69.76, 74.40) | 64.970.98 [-1pt](63.15, 67.04) | 67.341.09 [-1pt](65.15, 69.47) | 72.461.51 [-1pt](69.55, 75.44) | 86.341.14 [-1pt](84.01, 88.57) | 85.441.20 [-1pt](83.13, 87.67) |
| Closed-ended | ||||||
| Accuracy | 77.592.06 [-1pt](73.55, 81.48) | 75.472.16 [-1pt](71.09, 79.76) | 77.912.05 [-1pt](73.97, 81.74) | 74.022.16 [-1pt](69.74, 78.20) | 86.231.67 [-1pt](82.82, 89.49) | 86.031.74 [-1pt](82.56, 89.23) |
| F1 Score | 74.092.09 [-1pt](70.23, 78.03) | 63.402.39 [-1pt](58.69, 68.05) | 66.852.30 [-1pt](62.29, 71.46) | 74.022.16 [-1pt](69.74, 78.20) | 86.231.67 [-1pt](82.82, 89.49) | 86.031.74 [-1pt](82.56, 89.23) |
| Overall | ||||||
| Exact Match | 37.031.32 [-1pt](34.39, 39.75) | 25.301.28 [-1pt](22.66, 27.83) | 28.541.30 [-1pt](25.94, 30.98) | 65.671.38 [-1pt](62.94, 68.35) | 84.401.05 [-1pt](82.32, 86.36) | 83.941.13 [-1pt](81.61, 86.06) |
| F1 Score | 39.901.30 [-1pt](37.25, 42.57) | 27.421.25 [-1pt](24.85, 29.89) | 31.851.26 [-1pt](29.29, 34.27) | 66.581.35 [-1pt](63.85, 69.20) | 84.841.05 [-1pt](82.74, 86.83) | 84.201.14 [-1pt](81.86, 86.29) |
| GPT Score | 74.171.08 [-1pt](72.17, 76.31) | 69.091.02 [-1pt](67.09, 71.12) | 71.481.02 [-1pt](69.45, 73.42) | 73.071.23 [-1pt](70.71, 75.48) | 86.300.95 [-1pt](84.43, 88.15) | 85.671.02 [-1pt](83.51, 87.52) |
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | MedGemma | Gemma3 | MedGemma | Gemma3 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | LLaVA-Med | LLaVA-1.5 | LLaVA-Med | LLaVA-1.5 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | VILA-M3 | VILA | VILA-M3 | VILA |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |
| Metric | InternVL3 | InternVL3 |
| Open-ended | ||
| BLEU-1 | ||
| ROUGE-L | ||
| Exact Match | ||
| Accuracy | ||
| F1 Score | ||
| GPT Score | ||
| Closed-ended | ||
| Accuracy | ||
| F1 Score | ||
| Overall | ||
| Exact Match | ||
| F1 Score | ||
| GPT Score | ||
| Off-the-shelf | Fine-tuned (ML) | |||||
| Metric | Lingshu | Qwen25-VL | Qwen2-VL | Lingshu | Qwen25-VL | Qwen2-VL |
| Open-ended | ||||||
| BLEU-1 | 7.021.15 [-1pt](5.00, 9.37) | 1.370.10 [-1pt](1.18, 1.57) | 2.650.33 [-1pt](2.05, 3.36) | 29.253.14 [-1pt](22.97, 35.63) | 37.823.31 [-1pt](31.29, 44.65) | 37.303.12 [-1pt](31.38, 43.70) |
| ROUGE-L | 9.691.22 [-1pt](7.50, 12.14) | 2.540.18 [-1pt](2.20, 2.88) | 4.650.50 [-1pt](3.71, 5.71) | 31.723.19 [-1pt](25.24, 38.19) | 40.523.32 [-1pt](33.99, 47.25) | 39.833.14 [-1pt](33.67, 46.34) |
| Exact Match | 7.451.25 [-1pt](5.18, 10.02) | 1.370.10 [-1pt](1.18, 1.57) | 2.650.33 [-1pt](2.05, 3.36) | 33.893.39 [-1pt](27.03, 40.66) | 42.923.43 [-1pt](36.27, 49.62) | 41.733.24 [-1pt](35.34, 48.16) |
| Accuracy | 39.553.54 [-1pt](32.98, 46.23) | 45.543.58 [-1pt](38.61, 52.28) | 37.053.49 [-1pt](30.36, 43.66) | 25.483.28 [-1pt](19.17, 32.16) | 32.403.45 [-1pt](25.52, 39.20) | 33.433.45 [-1pt](27.14, 40.88) |
| F1 Score | 11.761.30 [-1pt](9.41, 14.43) | 5.060.38 [-1pt](4.31, 5.80) | 6.430.58 [-1pt](5.35, 7.60) | 31.753.21 [-1pt](25.34, 38.28) | 40.593.32 [-1pt](34.05, 47.35) | 40.053.14 [-1pt](33.92, 46.51) |
| GPT Score | 63.182.25 [-1pt](58.72, 67.48) | 61.971.81 [-1pt](58.40, 65.48) | 59.951.99 [-1pt](55.90, 63.59) | 54.812.56 [-1pt](49.78, 59.71) | 59.102.64 [-1pt](53.72, 64.04) | 56.032.59 [-1pt](50.90, 61.17) |
| Closed-ended | ||||||
| Accuracy | 70.522.98 [-1pt](64.62, 76.08) | 70.752.94 [-1pt](64.83, 76.33) | 75.412.74 [-1pt](70.19, 80.80) | 65.673.00 [-1pt](59.76, 71.16) | 80.842.46 [-1pt](76.03, 85.54) | 77.262.61 [-1pt](71.98, 82.38) |
| F1 Score | 70.522.98 [-1pt](64.62, 76.08) | 68.642.97 [-1pt](62.92, 74.40) | 75.412.74 [-1pt](70.19, 80.80) | 65.673.00 [-1pt](59.76, 71.16) | 80.842.46 [-1pt](76.03, 85.54) | 77.262.61 [-1pt](71.98, 82.38) |
| Overall | ||||||
| Exact Match | 42.512.30 [-1pt](37.67, 47.14) | 38.672.27 [-1pt](34.15, 43.17) | 43.092.23 [-1pt](38.52, 47.48) | 51.562.36 [-1pt](47.05, 56.20) | 64.002.20 [-1pt](59.62, 68.33) | 61.482.14 [-1pt](57.36, 65.60) |
| F1 Score | 44.422.25 [-1pt](39.86, 48.89) | 40.402.22 [-1pt](35.97, 44.81) | 44.772.18 [-1pt](40.39, 49.07) | 50.612.34 [-1pt](46.21, 55.21) | 62.972.18 [-1pt](58.59, 67.29) | 60.732.13 [-1pt](56.56, 64.83) |
| GPT Score | 67.261.89 [-1pt](63.61, 70.73) | 66.851.84 [-1pt](63.10, 70.27) | 68.551.81 [-1pt](65.00, 71.93) | 60.852.04 [-1pt](56.87, 64.96) | 71.191.87 [-1pt](67.55, 74.95) | 67.831.84 [-1pt](64.26, 71.50) |
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | MedGemma | Gemma3 | MedGemma | Gemma3 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | LLaVA-Med | LLaVA-1.5 | LLaVA-Med | LLaVA-1.5 |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |||
| Metric | VILA-M3 | VILA | VILA-M3 | VILA |
| Open-ended | ||||
| BLEU-1 | ||||
| ROUGE-L | ||||
| Exact Match | ||||
| Accuracy | ||||
| F1 Score | ||||
| GPT Score | ||||
| Closed-ended | ||||
| Accuracy | ||||
| F1 Score | ||||
| Overall | ||||
| Exact Match | ||||
| F1 Score | ||||
| GPT Score | ||||
| Off-the-shelf | Fine-tuned (ML) | |
| Metric | InternVL3 | InternVL3 |
| Open-ended | ||
| BLEU-1 | ||
| ROUGE-L | ||
| Exact Match | ||
| Accuracy | ||
| F1 Score | ||
| GPT Score | ||
| Closed-ended | ||
| Accuracy | ||
| F1 Score | ||
| Overall | ||
| Exact Match | ||
| F1 Score | ||
| GPT Score | ||